# linecolor=white, backgroundcolor=gray!20

### v1

@@ Abstract
Proteins play crucial roles in biological processes, with therapeutic peptides emerging as promising pharmaceutical agents. They allow new possibilities to leverage target binding sites that were previously undruggable. While deep learning (DL) has advanced peptide discovery, generating D-proteins composed of D-amino acids remains challenging due to the scarcity of natural examples. This paper proposes D-Flow, a full-atom flow-based framework for de novo D-peptide design. D-Flow is conditioned on receptor binding and utilizes a comprehensive representation of peptide structure, incorporating backbone frames, side-chain angles, and discrete amino acid types. A mirror-image algorithm is implemented to address the lack of training data for D-proteins, which converts L-receptors’ chirality. Furthermore, we enhance D-Flow’s capacity by integrating large protein language models (PLMs) with structural awareness through a lightweight structural adapter. A two-stage training pipeline and a controlling toolkit also enable D-Flow to transition from general protein design to targeted binder design while preserving pretraining knowledge. Extensive experimental results on the PepMerge benchmark demonstrate D-Flow’s effectiveness, particularly in developing peptides with entire D-residues. This approach represents a significant advancement in computational D-peptide design, offering unique opportunities for bioorthogonal and stable molecular tools and diagnostics. The code is available in https://github.com/smiles724/PeptideDesign
@@ 1 Introduction
Proteins are the building blocks of life and make essential contributions to nearly all fundamental biological processes in the cell. They fold into specific 3D conformations to perform distinct functionalities [27, 37]. Remarkably, therapeutic peptides, compromising a limited number of well-ordered residues, are single-chain proteins and an irregular class of pharmaceutical agents [69]. Peptide drugs occupy a unique chemical and pharmacological space between small and large molecules [42]. Due to this specialty, increasing efforts have adopted deep learning (DL) algorithms to facilitate their discovery [36, 41]. Notably, unbound peptide chains typically have high free energy and entropy, resulting in unstable conformations. In contrast, they can elicit pharmacological effects upon binding to specific receptors, forming stable complexes where both the receptor and ligand adopt equilibrium structures [31].
@@
Diffusion or flow models [44, 30, 60], a recent generative family, are revolutionizing the field of image generation [17], text-to-image generation [50], image segmentation [7], and also finding rapid applications in other domains such as video generation [29], and 3D generation [59]. Notable initiatives [71, 75] have explored the potential of diffusion or flow models in the context of drug discovery. These models have been applied to various protein representations, including carbon-alpha only structures [67], torsion angles [78], and the SE(3) backbone frame representation [80], and diverse scenes such as molecular design [28], antibody engineering [46, 48], de novo protein design [79], and peptide discovery [55] as well.
@@
Despite emerging interest in developing 3D diffusion or flow to peptides, little attention has been paid to D-protein generation. It is worth mentioning that though every cell in the human body contains proteins, their cornerstone, 20 categories of amino acids, can exist in two stereoisomers: L (levo) and D (dextro), which are non-superimposable mirror images of each other [38], and this chirality characteristic is determined by the orientation of the functional groups around the alpha carbon [24]. D-proteins are protein molecules whose polypeptide chains consist of D-amino acids and the achiral amino acid glycine. They can form specific heterochiral protein-protein interactions (PPIs) with natural L-protein targets and possess remarkable potential as therapeutics, and diagnostics because of their high bioorthogonality and stability [63]. However, in nature, the vast majority of peptides and proteins are made from L-amino acids, which poses a great challenge for DL mechanisms, as no training data is accessible.
@@
To overcome this obstacle, we introduce a mirror-imaged multi-modality model, named D-Flow (see Fig. 1), to produce D-peptides. D-Flow characterizes the peptide sequence and structure using rigid backbone frames within the SE(3) manifold, the side-chain angles on the high-dimensional torus, and the discrete amino acid types on the categorical space [10]. Each modality has an analytical flow, and they jointly capture the distribution of full-atom peptide structure conditioned on the receptor. During the inference phase, a novel mirror-image algorithm is adopted to transform the L-receptor and convert its natural chirality. Moreover, as datasets containing complete 3D structures of peptide complexes are orders of magnitude smaller than sequence-only databases [77], the scarcity of high-quality peptide-receptor pairs inevitably constrains the performance of DL approaches. As a resolution, we propose to empower the flow model with large protein language models (PLMs) and a two-stage training pipeline. Specifically, a structural surgery is implemented on PLMs using a lightweight structural adapter [84]. Consequently, PLMs are endowed with structural awareness and can facilitate the regression of residue categories given intermediate-state conformation information. Secondly, D-Flow embraces a controlling toolkit [82] to achieve awareness of target proteins and transfer from the general protein design task to the binder design task with the minimum loss of the pretraining knowledge. Comprehensive experiments on the PepMerge benchmark [41] verify the effectiveness of D-Flow, particularly, for its breakthrough in generating pure D-peptides. Besides, it also outpasses existing DL baselines in generating L-peptides by a large margin. Our study opens up avenues for systematically exploring the DL-based mirror-image protein universe, paving the way for a wide range of design applications targeting L-proteins.
Figure 1: Illustration of D-Flow.
2 Related Works
2.1 Diffusion and Flow for Protein Design
Generative models have displayed incredible promise in designing novel protein structures for custom functions, with much of this progress driven by advancements in diffusion models [60, 30, 17]. Among distinct protein representations, the frame representation has achieved state-of-the-art performance in tasks like de novo backbone design and motif scaffolds [67], as exemplified by RFdiffusion [71]. Recently, flow methods [44, 64] appear as an alternative to diffusion models, offering a deterministic approach by eliminating stochasticity from the sampling process. They have been extended to Riemannian manifolds [44, 12] and possess an attractive connection to optimal transport (OT). The linear interpolation schedule of flow results in more direct sampling trajectories, reducing the number of integration steps needed [44]. In the computer vision field, flow matching (FM) has already demonstrated comparable results to diffusion models with more computational efficiency [53].
The past few years have witnessed a growth in applying diffusion and flow models [46, 76, 48, 35] for antibody design, a special protein family in the immune system to resist antigens, mainly focusing on inpainting complementarity-determining regions (CDRs) at the interface between the antigen and the framework. Lately, non-trivial efforts have been put into developing DL algorithms for peptides. Li et al. [41] encompasses the dynamics of side-chain angles. Contrarily, Lin et al. [42] relies on RDE [47] to complete the solution to full-atom design. Kong et al. [36] presents a full-atom diffusion on the latent geometric space learned by VAE. Nevertheless, none considers the extension of sequence models (e.g., PLMs) and multi-stage training paradigms to alleviate the shortage of available binding structures.
2.2 D-peptide Technology
Proteins are built from chiral molecules, specifically amino acids that exist in two mirror-image forms: L (Levorotatory) and D (Dextrorotatory) [16]. Natural ribosomes exclusively use L-amino acids to synthesize proteins, resulting in L-framework proteins throughout nature. While engineered ribosomes [26], post-translational modification systems [34], and non-ribosomal peptide synthetases [51] can incorporate some D-amino acids into L-protein chains, proteins made entirely of D-amino acids have never been found in nature [38] and must be chemically synthesized in the laboratory. Despite the synthetic challenges, D-proteins are valuable research tools. They fold into mirror images of their L-counterparts and offer unique opportunities for studying fundamental protein mechanisms, developing stable molecular binding agents, and even exploring the possibility of mirror-image biological systems [49, 19]. To identify D-proteins capable of binding to a target L-protein, mirror-image peptide phage display methods have been developed [58, 11, 85, 9, 54], which involves screening a phage library of L-peptides with a target D-protein. However, it remains challenging to precisely target a specific surface region of the target protein and confirm the presence of valid binders within the initial random library [63]. More importantly, no preceding works have explored the possibility of DL co-design methods [46, 80] for D-protein generation, not to mention the incorporation and verification of this mirror-image technique.
3 Preliminary and Background
Proteins.
A protein is a biomolecule consisting of $n$ amino acid residues, each defined by its type, backbone frame, and side-chain torsion angles [22]. The type of the $i$-th residue, denoted by $a^{i}\in\mathcal{A}$, is determined by its side-chain R group, where $\mathcal{A}=\{1\ldots 20\}$. The rigid frame of each residue is constructed from the coordinates of four backbone heavy atoms $\mathrm{N}$, $\mathrm{C}\alpha$, $\mathrm{C}$, and $\mathrm{O}$, with $\mathrm{C}\alpha$ positioned at the origin. This frame is represented by a position vector $\boldsymbol{x}^{i}\in\mathbb{R}^{3}$ and a rotation matrix $O^{i}\in\mathrm{SO}(3)$ [32]. Additionally, seven torsion angles $\boldsymbol{\chi}^{i}\in[0,2\pi)^{7}$ are considered and three of them belong to the backbone, where $\boldsymbol{\chi}^{i}[0]$ is the angle around $\mathrm{C}-\mathrm{N}$ bond, $\boldsymbol{\chi}^{i}[1]$ is the angle around $\mathrm{N}-\mathrm{C}\alpha$ bond, and $\boldsymbol{\chi}^{i}[2]$ affects the position of the oxygen atom. Unlike the backbone, the side-chain conformation is more flexible, involving up to four rotatable torsion angles between side-chain atoms, denoted by $\boldsymbol{\chi}^{i}[3:]\in[0,2\pi)^{4}$.
Flow on Riemannian Manifolds.
Flow matching (FM) [44, 1, 2, 64] is a simulation-free method for learning a continuous normalizing flow (CNF). On a manifold $\mathcal{M}$ with a Riemannian metric $g$, the CNF $\Phi:\mathcal{M}\rightarrow\mathcal{M}$ is defined by a one-parameter diffeomorphism that integrates along a time-dependent vector field $u_{t}\in\mathcal{U}$, where $u_{t}(x)\in\mathcal{T}_{x}\mathcal{M}$ falls at the tangent space of the manifold at $x\in\mathcal{M}$.
With an initial condition of $\phi_{0}(x)=x$, the time-dependent flow $\phi_{t}:\mathcal{M}\rightarrow\mathcal{M}$ and the final diffeomorphismis are attained by setting $\Phi(x)=\phi_{1}(x)$ and solving the ordinary differential equation (ODE) $\frac{\mathrm{d}\phi_{t}}{\mathrm{d}t}({x})=u_{t}(\phi_{t}(x))$ on $\mathcal{M}$ over $t\in[0,1]$. Remarkably, $\phi_{t}(x)$ transports the point $x$ along the vector field $u_{t}(x)$ from time 00 up to time $t$, satisyfing another ODE: $\mathrm{d}x=u_{t}(x)\mathrm{d}t$.
Denote $\mathbb{P}(\mathcal{M})$ as the space of probability distributions on $\mathcal{M}$. Then the flow can transform a simple prior density $p_{0}\in\mathbb{P}(\mathcal{M})$ towards the data distribution $p_{1}\in\mathbb{P}(\mathcal{M})$ via the push-forward equation $p_{t}=[\phi_{t}]_{\#}p_{0}$, and its density is $p_{t}(x)=[\phi_{t}]_{\#}p_{0}(x)=p_{0}\left(\phi_{t}^{-1}(x)\right)\textrm{det}\left[\frac{\partial\phi_{t}^{-1}}{\partial x}(x)\right]$. This sequence of probability distributions $\{p_{t}:t\in[0,1]\}$ is referred to the probability path, which interpolates $\mathbb{P}(\mathcal{M})$ between $p_{0}$ and $p_{1}$.
[…]
@@ 4.1 Flow for Full-atom Protein Generation
Discrete Flow for 1D Amino Acid Sequence.
Though flow models succeeds in continuous spatial signals like images [30, 57], they falter when applied to discrete sequential data. Two recent lines adapt diffusion or flow to the discrete setting: one embeds the discrete data in continuous space and adopts the continuous diffusion or flow [18, 41, 70], and the other designs the transformation process immediately over categorical state spaces [6, 10, 25]. In this work, we investigate both approaches and observe a significant advantage of discrete flow over the continuous version.
Following Campbell et al. [10] and Gat et al. [25], we define FM in a discrete setting and consider a Continuous-Time discrete Markov Chain (CTMC) paradigm. The categorical variable $a^{i}_{t}$ jumps between states in the amino acid type space $\mathcal{A}$ depending on a continuous time $t\in[0,1]$. $u_{t}$, called the probability velocity as reminiscent of the velocity field in CNF, is defined as the rate of probability change of the current sample $a_{t}^{i}$ in each of its $\|\mathcal{A}\|=20$ categories. Thus, each category of the residue type $a^{i}_{t}\sim p_{t}$ is updated independently with the Euler step as $a_{t+\frac{1}{N}}^{i}\sim\delta_{a_{t}^{i}}(\cdot)+\frac{1}{N}u_{t}\left(a^{i},\cdot\right)$. For a sufficiently large number of timesteps $N\rightarrow\infty$ and any potential state $z\in\mathcal{A}$, the probability velocity $u_{t}$ is required to satisfy
@@
[…]
@@
where $p_{1\mid t}\left(a^{i}\mid z\right)=\sum_{a_{0}^{i},a_{1}^{i}}\delta_{a_{1}^{i}}\left(a^{i}\right)p_{t}\left(a_{0}^{i},a_{1}^{i}\mid z\right)$ is used as the notation for the probability denoiser. The training goal is to learn this denoiser $p_{1\mid t}$ by minimizing the cross-entropy (CE) loss: {mdframed}
$\begin{split}\mathcal{L}_{\text{aa}}(\theta)=&\mathbb{E}_{i\in[n],t\sim\mathcal{U}(0,1),p_{0}\left(a^{i}\right),p_{1}\left(a^{i}\right),p_{t}\left(a^{i}|a^{i}_{0},a^{i}_{1}\right)}\\ &\left[\log p_{1\mid t}\left(a^{i}_{1}\mid a^{i}_{t}\right)\right].\end{split}$ (4)
@@
Multi-modality FM for 3D Protein Structures.
Apart from the sequential amino acid type, we follow common practice [79, 42, 41, 10] and construct different probability flows containing Euclidean, spherical, and toric FMs for positions $\boldsymbol{x}\in\mathbb{R}^{n\times 3}$, orientations $O\subseteq\mathrm{SO}(3)$, and torsion angles $\boldsymbol{\chi}\in[0,2\pi)^{n\times 7}$, respectively. To be specific, a vanilla Gaussian FM on Euclidean manifolds with the prior $\mathcal{N}\left(0,\boldsymbol{I}_{3}\right)$ is leveraged to generate $\boldsymbol{x}^{i}$. As for 3D rotation group $\mathrm{SO}(3)$, it is also a smooth Riemannian manifold with its tangent space $\mathfrak{so}(3)$ forming a Lie algebra consisting of skew-symmetric matrice. We establish flows based on the geodesics in the context of $\mathrm{SO}(3)$ [39] and select the uniform distribution over $\mathrm{SO}(3)$ as the prior $p\left(O_{0}^{i}\right)$. For torsion angles $\boldsymbol{\chi}^{i}\in[0,2\pi)^{7}$, each of them can be represented as a point on the unit circle $\mathbb{S}^{1}$. Thus, $\boldsymbol{\chi}^{i}$ lies on the 7-dimensional hypertorus $\mathbb{T}^{7}=(\mathbb{S}^{1})^{7}$ as the Cartesian product of all seven unit circles. This flat torus $\mathbb{T}^{7}$ can be viewed as the quotient space $(\mathbb{R}^{7}\mod 2\pi\mathbb{Z})^{7}$ that inherits the Riemannian metric from Euclidean space, where the uniform distribution on $[0,2\pi)^{7}$ is utilized as the prior. The conditional flows are prescribed as:
@@
$\phi_{t}^{\text{ang}}(\boldsymbol{\chi}_{0}^{i},\boldsymbol{\chi}_{1}^{i})$ $=\left[t\boldsymbol{\chi}_{1}^{i}+(1-t)\boldsymbol{\chi}_{0}^{i}\right]\mod 2\pi,$ (8)
$\quad\boldsymbol{\chi}_{0}^{i}\sim\mathcal{U}\left([0,2\pi)^{7}\right),$ (9)
@@
Ultimately, time-dependent and equivariant neural networks $v^{\text{pos}}(\cdot)$, $v^{\text{ori}}(\cdot)$, and $v^{\text{ang}}(\cdot)$, which share the same backbone architecture as $v^{\text{aa}}(\cdot)$ but have different head predictors, are employed to approximate the conditional gradient fields $u_{t}^{\text{pos}}$, $u_{t}^{\text{ori}}$, and $u_{t}^{\text{ang}}$. The corresponding FM objectives are given below, separately: {mdframed}
$\begin{split}\mathcal{L}_{\text{pos}}(\theta)=&\mathbb{E}_{i\in[n],t\sim\mathcal{U}(0,1),p\left(\boldsymbol{x}_{1}^{i}\right),p\left(\boldsymbol{x}_{0}^{i}\right),p\left(\boldsymbol{x}_{t}^{i}|\boldsymbol{x}_{0}^{i},\boldsymbol{x}_{1}^{i}\right)}\\ &\left\|v^{\text{pos}}\left(\boldsymbol{x}_{t}^{i},t\right)-\left(\boldsymbol{x}_{1}^{i}-\boldsymbol{x}_{0}^{i}\right)\right\|_{2}^{2},\\ \end{split}$ (13)
$\begin{split}\mathcal{L}_{\text{ori}}(\theta)=&\mathbb{E}_{i\in[n],t\sim\mathcal{U}(0,1),p\left(O_{1}^{i}\right),p\left(O_{0}^{i}\right),p\left(O_{t}^{i}|O_{0}^{i},O_{1}^{i}\right)}\\ &\left\|v^{\text{ori}}\left(O_{t}^{i},t\right)-\frac{\log_{O_{t}^{i}}\left(O_{1}^{i}\right)}{1-t}\right\|_{\mathrm{SO}(3)}^{2},\\ \end{split}$ (14)
$\begin{split}\mathcal{L}_{\text{ang}}(\theta)=&\mathbb{E}_{i\in[n],t\sim\mathcal{U}(0,1),p\left(\boldsymbol{\chi}_{1}^{i}\right),p\left(\boldsymbol{\chi}_{0}^{i}\right),p\left(\boldsymbol{\chi}_{t}^{i}|\boldsymbol{\chi}_{0}^{i},\boldsymbol{\chi}_{1}^{i}\right)}\\ &\left\|v^{\text{ang}}\left(\boldsymbol{\chi}_{t}^{i},t\right)-(\boldsymbol{\chi}_{1}^{i}-\boldsymbol{\chi}_{0}^{i})\right\|_{2}^{2}.\end{split}$ (15)
We summarize losses of all different modalities to acquire the overall FM objective as $\mathcal{L}_{\mathrm{CFM}}=\lambda_{\text{pos}}\mathcal{L}_{\text{pos}}+\lambda_{\mathrm{ori}}\mathcal{L}_{\text{ori}}+\lambda_{\text{ang}}\mathcal{L}_{\text{ang}}+\lambda_{\text{aa}}\mathcal{L}_{\text{aa}}$, where $\lambda_{*}$ are the hyperparameters to control the impact of different loss components. Two additional losses are also imposed concerning the backbone atoms and distance matrix [79].
@@
$\boldsymbol{x}_{t+\frac{1}{N}}^{i}$ $=\boldsymbol{x}_{t}^{i}+\frac{1}{N}v^{\text{pos}}\left(\boldsymbol{x}_{t}^{i},t\right),$ (16)
$O_{t+\frac{1}{N}}^{i}$ $=\exp_{O_{t}^{i}}\left(\frac{1}{N}v^{\text{ori}}\left(O_{t}^{i},t\right)\right),$ (17)
$\boldsymbol{\chi}_{t+\frac{1}{N}}^{i}$ $=\left[\boldsymbol{\chi}_{t}+\frac{1}{N}v^{\text{ang}}\left(\boldsymbol{\chi}_{t}^{i},t\right)\right]\mod 2\pi.$ (18)

### v3

# D-Flow: Multi-modality Flow Matching for D-peptide Design

@@ Abstract
Proteins play crucial roles in biological processes, with therapeutic peptides emerging as promising pharmaceutical agents. They allow for new possibilities to leverage target binding sites that were previously undruggable. Although deep learning has advanced peptide discovery, generating D-proteins composed of D-amino acids remains challenging because of the scarcity of natural examples. This paper proposes D-Flow, a full-atom flow-based framework for de novo D-peptide design. D-Flow is conditioned on receptor binding and utilizes a comprehensive representation of peptide structure, incorporating backbone frames, side-chain angles, and discrete amino acid types. A mirror-image algorithm is implemented to address the lack of training data for D-proteins, which convert the chirality of L-receptors. Furthermore, we enhance D-Flow’s capacity by integrating large protein language models with structural awareness through a lightweight structural adapter. A two-stage training pipeline and a controlling toolkit also enable D-Flow to transition from a general protein design to a targeted binder design while preserving pre-training knowledge. Extensive experimental results on the PepMerge benchmark demonstrate D-Flow’s effectiveness, particularly in developing peptides with entire D-residues. This approach represents a significant advancement in computational D-peptide design, offering unique opportunities for bioorthogonal and stable molecular tools and diagnostics.
@@ 1 Introduction
Proteins are the building blocks of life and make essential contributions to nearly all fundamental biological processes in the cell. They fold into 3D conformations to perform distinct functionalities. Therapeutic peptides, comprising a limited number of well-ordered residues, are single-chain proteins and an irregular class of pharmaceutical agents [51, 56, 57]. Peptide drugs occupy a unique chemical and pharmacological space between small and large molecules [27]. Due to this specialty, increasing efforts have adopted deep learning (DL) algorithms to facilitate their discovery [26, 63]. Notably, unbound peptide chains typically have high free energy and entropy, resulting in unstable conformations. In contrast, they can elicit pharmacological effects upon binding to specific receptors, forming stable complexes where both the receptor and ligand adopt equilibrium structures.
@@
Diffusion or flow models [29, 58], a recent generative family, are revolutionizing several domains. Notable initiatives have explored the potential of diffusion or flow models in the context of drug discovery. These models have been applied to various protein representations, including carbon-alpha only structures [49], torsion angles [65], and the SE(3) backbone frame representation [67], and diverse scenes such as molecular design [16], antibody engineering [33], de novo protein design [66], and peptide discovery [38] as well. For a more detailed description, see Appendix A.
@@
Despite emerging interest in developing 3D diffusion or flow to peptides, little attention has been paid to D-protein generation in machine learning (ML). It is worth mentioning that though every cell in the human body contains proteins, their cornerstone, 20 categories of amino acids, can exist in two stereoisomers: L (levo) and D (dextro), which are non-superimposable mirror images of each other [24], and this chirality characteristic is determined by the orientation of the functional groups around the alpha carbon. D-proteins are protein molecules whose polypeptide chains consist of D-amino acids and the achiral amino acid glycine. They can form specific heterochiral protein-protein interactions (PPIs) with natural L-protein targets and possess remarkable potential as therapeutics, and diagnostics because of their high bioorthogonality and stability [45]. However, in nature, the vast majority of peptides and proteins are made from L-amino acids, which poses a great challenge for DL mechanisms, as no training data is accessible.
Figure 1: Illustrative overview of our D-Flow framework.
@@
To overcome this obstacle, we introduce a mirror-imaged multi-modality model, named D-Flow (see Fig. 1), to produce D-peptides. D-Flow characterizes the peptide sequence and structure using rigid backbone frames within the SE(3) manifold, the side-chain angles on the high-dimensional torus, and the discrete amino acid types on the categorical space [5]. Each modality has an analytical flow, and they jointly capture the distribution of full-atom peptide structure conditioned on the receptor. During the inference phase, a novel mirror-image algorithm is adopted to transform the L-receptor and convert its natural chirality. Moreover, as datasets containing complete 3D structures of peptide complexes are orders of magnitude smaller than sequence-only databases [62], the scarcity of high-quality peptide-receptor pairs inevitably constrains the performance of DL approaches. As a resolution, we propose to empower the flow model with large protein language models (PLMs) and a two-stage training pipeline. Specifically, a structural surgery is implemented on PLMs using a lightweight structural adapter [71]. Consequently, PLMs are endowed with structural awareness and can facilitate the regression of residue categories given intermediate-state conformation information. Secondly, D-Flow embraces a controlling toolkit [69] to achieve awareness of target proteins and transfer from the general protein design task to the binder design task with the minimum loss of the pretraining knowledge. Comprehensive experiments on the PepMerge benchmark [26] verify the effectiveness of D-Flow, particularly for its breakthrough in generating pure D-peptides. Besides, it outpaces existing DL baselines in generating L-peptides by a large margin. Our study opens up avenues for systematically exploring the DL-based mirror-image protein universe, paving the way for various design applications targeting L-proteins.
@@ 2.1 Flow for Full-atom Protein Generation
Discrete Flow for 1D Amino Acid Sequence. We provide the preliminary and background of Proteins and Flow on Riemannian Manifolds in Appendix B. Although flow models succeed in continuous spatial signals like images [40], they falter when applied to discrete sequential data. Two recent lines adapt diffusion or flow to the discrete setting: one embeds the discrete data in continuous space and adopts the continuous diffusion or flow [26], and the other designs the transformation process immediately over categorical state spaces [5, 14]. In this work, we investigate both approaches and observe a significant advantage of discrete flow over the continuous version. Following Campbell et al. [5] and Gat et al. [14], we define FM in a discrete setting and consider a Continuous-Time discrete Markov Chain (CTMC) paradigm. The categorical variable $a^{i}_{t}$ jumps between states in the amino acid type space $\mathcal{A}$ depending on a continuous time $t\in[0,1]$. $u_{t}$, called the probability velocity as reminiscent of the velocity field in CNF, is defined as the rate of probability change of the current sample $a_{t}^{i}$ in each of its $\|\mathcal{A}\|=20$ categories. Thus, each category of the residue type $a^{i}_{t}\sim p_{t}$ is updated independently with the Euler step as $a_{t+\frac{1}{N}}^{i}\sim\delta_{a_{t}^{i}}(\cdot)+\frac{1}{N}u_{t}\left(a^{i},\cdot\right)$. For a sufficiently large number of timesteps $N\rightarrow\infty$ and any potential state $z\in\mathcal{A}$, $u_{t}$ satisfies
@@
$\text{ for }\forall i\in[n]\text{ and }a^{i}\neq z$. Then, the unconditional probability velocity $u_{t}$ can be constructed as the marginalization of the conditional one $u_{t}\left(a^{i},z\mid a^{i}_{0},a^{i}_{1}\right)$. As a consequence, we attain the marginal velocity written as $u_{t}\left(a^{i},z\right)=\sum_{a^{i}_{0},a^{i}_{1}\in\mathcal{A}}u_{t}\left(a^{i},z\mid a^{i}_{0},a^{i}_{1}\right)p_{t}\left(a^{i}_{0},a^{i}_{1}\mid z\right)$, which generates the probability path $p_{t}\left(a^{i}\right)$. Generally, the conditional probability paths can be represented as a convex sum of $m$ conditional probabilities, namely, $p_{t}\left(a^{i}\mid a^{i}_{0},a^{i}_{1}\right)=\sum_{j=1}^{k}\kappa_{t}^{j}p^{j}(a^{i}\mid a^{i}_{0},a^{i}_{1})$, where the schedulers $\kappa_{t}^{j}$ are collectively non-negative and satisfy $\sum_{j}\kappa_{t}^{j}=1$. A simple yet useful instance of conditional probability paths is $p_{t}\left(a^{i}\mid a^{i}_{0},a^{i}_{1}\right)=\kappa_{t}\delta_{a^{i}_{1}}\left(a^{i}\right)+(1-\kappa_{t})\delta_{a^{i}_{0}}\left(a^{i}\right)$, where the scheduler $\kappa_{t}$ satisfies $\kappa_{0}=0$, $\kappa_{1}=1$ and monotonically increases in $t$ (i.e., $\dot{\kappa}_{t}\geq 0$). This results in $p_{0}(a^{i}|a^{i}_{0},a^{i}_{1})=\delta_{a^{i}_{0}}(a^{i})$ and $p_{1}(a^{i}|a^{i}_{0},a^{i}_{1})=\delta_{a^{i}_{1}}(a^{i})$. Subsequently, we get:
@@
where $p_{1\mid t}\left(a^{i}\mid z\right)=\sum_{a_{0}^{i},a_{1}^{i}}\delta_{a_{1}^{i}}\left(a^{i}\right)p_{t}\left(a_{0}^{i},a_{1}^{i}\mid z\right)$ is used as the notation for the probability denoiser. The training goal is to learn $p_{1\mid t}$ by minimizing the cross-entropy (CE) loss:
$\begin{split}\mathcal{L}_{\text{aa}}(\theta)=&\mathbb{E}_{i\in[n],t\sim\mathcal{U}(0,1),p_{0}\left(a^{i}\right),p_{1}\left(a^{i}\right),p_{t}\left(a^{i}|a^{i}_{0},a^{i}_{1}\right)}\left[\log p_{1\mid t}\left(a^{i}_{1}\mid a^{i}_{t}\right)\right].\end{split}$ (3)
@@
Multi-modality FM for 3D Protein Structures. Apart from the sequential amino acid type, we follow common practice [66, 27, 26, 5] and construct different probability flows containing Euclidean, spherical, and toric FMs for positions $\boldsymbol{x}\in\mathbb{R}^{n\times 3}$, orientations $O\subseteq\mathrm{SO}(3)$, and torsion angles $\boldsymbol{\chi}\in[0,2\pi)^{n\times 7}$, respectively. To be specific, a vanilla Gaussian FM on Euclidean manifolds with the prior $\mathcal{N}\left(0,\boldsymbol{I}_{3}\right)$ is leveraged to generate $\boldsymbol{x}^{i}$. As for the 3D rotation group $\mathrm{SO}(3)$, it is also a smooth Riemannian manifold with its tangent space $\mathfrak{so}(3)$ forming a Lie algebra consisting of skew-symmetric matrice. We establish flows based on the geodesics in the context of $\mathrm{SO}(3)$ and select the uniform distribution over $\mathrm{SO}(3)$ as the prior $p\left(O_{0}^{i}\right)$. For torsion angles $\boldsymbol{\chi}^{i}\in[0,2\pi)^{7}$, each can be represented as a point on the unit circle $\mathbb{S}^{1}$. Thus, $\boldsymbol{\chi}^{i}$ lies on the 7-dimensional hypertorus $\mathbb{T}^{7}=(\mathbb{S}^{1})^{7}$ as the Cartesian product of all seven unit circles. This flat torus $\mathbb{T}^{7}$ can be viewed as the quotient space $(\mathbb{R}^{7}\mod 2\pi\mathbb{Z})^{7}$ that inherits the Riemannian metric from Euclidean space, where the uniform distribution on $0,2\pi)^{7}$ is utilized as the prior. The conditional flows are prescribed as:
@@
$\phi_{t}^{\text{ang}}(\boldsymbol{\chi}_{0}^{i},\boldsymbol{\chi}_{1}^{i})$ $=\left[t\boldsymbol{\chi}_{1}^{i}+(1-t)\boldsymbol{\chi}_{0}^{i}\right]\mod 2\pi,\quad\boldsymbol{\chi}_{0}^{i}\sim\mathcal{U}\left([0,2\pi)^{7}\right),$ (7)
@@
Ultimately, time-dependent equivariant neural networks $v^{\text{pos}}(\cdot)$, $v^{\text{ori}}(\cdot)$, and $v^{\text{ang}}(\cdot)$, with the same backbone architecture as $v^{\text{aa}}(\cdot)$ but different head predictors, are employed to approximate the conditional gradient fields $u_{t}^{\text{pos}}$, $u_{t}^{\text{ori}}$, and $u_{t}^{\text{ang}}$. The FM objectives are:
$\mathcal{L}_{\text{pos}}(\theta)=\mathbb{E}_{i\in[n],t\sim\mathcal{U}(0,1),p\left(\boldsymbol{x}_{1}^{i}\right),p\left(\boldsymbol{x}_{0}^{i}\right),p\left(\boldsymbol{x}_{t}^{i}|\boldsymbol{x}_{0}^{i},\boldsymbol{x}_{1}^{i}\right)}\left\|v^{\text{pos}}\left(\boldsymbol{x}_{t}^{i},t\right)-\left(\boldsymbol{x}_{1}^{i}-\boldsymbol{x}_{0}^{i}\right)\right\|_{2}^{2},$ (11)
$\mathcal{L}_{\text{ori}}(\theta)=\mathbb{E}_{i\in[n],t\sim\mathcal{U}(0,1),p\left(O_{1}^{i}\right),p\left(O_{0}^{i}\right),p\left(O_{t}^{i}|O_{0}^{i},O_{1}^{i}\right)}\left\|v^{\text{ori}}\left(O_{t}^{i},t\right)-\frac{\log_{O_{t}^{i}}\left(O_{1}^{i}\right)}{1-t}\right\|_{\mathrm{SO}(3)}^{2},$ (12)
$\mathcal{L}_{\text{ang}}(\theta)=\mathbb{E}_{i\in[n],t\sim\mathcal{U}(0,1),p\left(\boldsymbol{\chi}_{1}^{i}\right),p\left(\boldsymbol{\chi}_{0}^{i}\right),p\left(\boldsymbol{\chi}_{t}^{i}|\boldsymbol{\chi}_{0}^{i},\boldsymbol{\chi}_{1}^{i}\right)}\left\|v^{\text{ang}}\left(\boldsymbol{\chi}_{t}^{i},t\right)-(\boldsymbol{\chi}_{1}^{i}-\boldsymbol{\chi}_{0}^{i})\right\|_{2}^{2}.$ (13)
The overall FM objective is the loss sum of all different modalities as $\mathcal{L}_{\mathrm{CFM}}=\lambda_{\text{pos}}\mathcal{L}_{\text{pos}}+\lambda_{\mathrm{ori}}\mathcal{L}_{\text{ori}}+\lambda_{\text{ang}}\mathcal{L}_{\text{ang}}+\lambda_{\text{aa}}\mathcal{L}_{\text{aa}}$, where $\lambda_{*}$ are balance weight hyperparameters. Two additional losses are also imposed concerning the backbone atoms and the distance matrix [66].
@@
$\boldsymbol{x}_{t+\frac{1}{N}}^{i}=\boldsymbol{x}_{t}^{i}+\frac{1}{N}v^{\text{pos}}\left(\boldsymbol{x}_{t}^{i},t\right),\,O_{t+\frac{1}{N}}^{i}=\exp_{O_{t}^{i}}\left(\frac{1}{N}v^{\text{ori}}\left(O_{t}^{i},t\right)\right),\,\boldsymbol{\chi}_{t+\frac{1}{N}}^{i}=\left[\boldsymbol{\chi}_{t}^{i}+\frac{1}{N}v^{\text{ang}}\left(\boldsymbol{\chi}_{t}^{i},t\right)\right]\text{mod}\,2\pi.$ (14)